Solve the following triangle for the remaining angles and sides. This triangle has 2 cases:
side a = 3
angle A = 10 degrees
side b = 4

Answer by stanbon(75887)

Solve the following triangle for the remaining angles and sides. This triangle has 2 cases:
side a = 3
angle A = 10 degrees
side b = 4
-----------------
Law of Sines::
sin(B)/4 = sin(10)/3
---
sin(B) = (4/3)sin(10) = 0.2315
B = arcsin(0.2315) = 13.39 degrees or B = 180-13.39 = 166.61 degrees
---
C = 180-(10 + 13.39) = 156.61 degrees OR C = 3.39 degrees
---
If C = 156.61 degrees,
c/sin(156.61) = 4/sin(10)
C = 0.3970*4/0.1736 = 9.145
---
If C = 3.39 degrees
C/sin(3.39) = 4/sin(10)
C = 1.36
---------------
Cheers,
Stan H.
---------------